A sample of 40 farmers comprises of 25 adopters and 15 non-adopters of organic fertilizer. Also 15 the adopters grow cowpea, while 7 of the non-adopters grow rice. Determine the probability that the first person interviewed was either a non-adopter or someone that grow cowpea


1
Expert's answer
2021-05-28T10:09:56-0400

"P=\frac{15}{40}+\frac{15}{40}=\frac{30}{40}=\frac{3}{4}"